It takes 49.0 J to raise the temperature of an 9.30 g piece of unknown metal from 13.0∘C to 25.0 ∘C. What is the specific heat for the metal?
Answer #1

heat taken(q) = m*s*DT

q = 49.0 j

m = mass of metal = 9.3 g

s = specific heat of metal = x j/g.c

DT = 25-13 = 12

49.0 = 9.3*x*12

x = 0.439

s = specific heat of metal = x = 0.439 j/g.c